# Investigating the Representation of Backchannels and Fillers in Fine-tuned Language Models

### v1

@@ 3.3 Experiment Set-Up
Our experiment workflow is shown in Figure 2. The chosen LMs are fine-tuned in advance based on the three different fine-tuning tasks [^5] Note that in the later stage, when we extract the embeddings of backchannels/fillers for the pre-trained LMs without fine-tuning, the pipeline works in the same way.. We selected 80% of the conversation data for fine-tuning and the remaining 20% for subsequent generative evaluation. For the evaluation of the representations of backchannels/fillers, we used all the conversation data. We use the backchannel ‘uh huh’ as an example to illustrate the workflow. In the initial stage, we pass through all input samples $\mid\chi\mid$ and filter out all utterances that contain ‘uh huh’. We then encode the corresponding utterances through the different hidden layers of a fine-tuned LM, obtaining the representation vector of ‘uh huh’ from the final hidden layer. Next, we check the dimensions of the representation matrix. Given that some of the backchannels/fillers can consist of more than one token (as ‘uh huh’ in our example), we then apply dimension reduction, simply taking the weighted average value of one dimension of the matrix so that the matrix can be levelled down to a vector representation (embedding). With the embeddings of the backchannels/fillers ‘uh huh’, we examine its representation via clustering. For BERT and GPT-2 models used in the fine-tuning task, the dimension of the hidden layer is 768 for English GPT-2 model and 1024 for the Japanese GPT-2 model. TurnGPT is based on the GPT-2, thus shares the same dimension. A special case is LLaMA-3 8B and Qwen-3 8B as their hidden layer dimension is 4096. For computational reasons, we reduce the dimensionality of the obtained embeddings from all of the models to 100 by using Principal Component Analysis (PCA) Abdi and Williams (2010). Technical details of the fine-tuning (e.g., GPU run time, usage of LoRA for parameter optimisation when fine-tuning LLaMA model Hu et al., 2022) can be found in Appendix C.
@@
Moreover, for extracting the embeddings of backchannels/fillers after fine-tuning, we have three different contextual settings: (i) no context information; (ii) one context information; (iii) full context information. In the first setting, when we encounter a backchannel/filler in an utterance, only the utterance containing that backchannel/filler is fed to obtain its embedding. In the second setting, we use the previous and subsequent utterances of a backchannel/filler to build the context before obtaining its embedding. For the third setting, we combine all previous utterances when we encounter an utterance containing a backchannel/filler, using this combination as the input to obtain the embedding of that backchannel/filler. The third setting is only applicable to the LLaMA-3 8B and Qwen-3 8B models since both have no input length limitation.
@@ 4 Analysis and Results
We chose the top 15 most frequent backchannels/fillers in our Japanese and English data to check how the representation of backchannels/fillers changes after fine-tuning. Among the 15 selected backchannels/fillers in each language, typical examples are: the ones indicating positive or negative feedback as the reference objects, and signals for turn-holding (in English, ‘yes’, ‘uh’, ‘yeah’ etc.; in Japanese, examples are ‘はい’(hai, ‘yes’ in English), ‘うん’(un, ‘yeah’), ‘ああ’(aa, ‘ah’) etc.).
We report the preliminary results using t-SNE visualization (Van der Maaten and Hinton, 2008). The embeddings of the 15 selected backchannels/fillers in each language are the input data. These embeddings are extracted from the last hidden layer of the LMs, before and after fine-tuning, as shown in Figure 2. We have selected t-SNE visualisations from the LLaMA-3 model with a single context setting to demonstrate how fine-tuning (in this instance, NTP) modifies the representation of backchannels/fillers within the language models (LMs). Figure 3a shows that for English data, when we obtain the embeddings of a selected backchannel/filler from the pre-trained LLaMA-3 model, the distinction between different backchannels/fillers are not clear enough (see Figure 3a). We believe that the large overlap of different data points is due to the fact that the pre-trained LMs have limited knowledge of the backchannel/filler and thus will assign random values to the encountered backchannel/filler. After fine-tuning, clearer distinction among the embeddings of different backchannels/fillers start to appear in the English data (see Figure 3b). The effect of fine-tuning for Japanese data is much clearer, after fine-tuning distinctions between different backchannels/fillers emerge as more colours appear in the t-SNE visualization (Figure 3c and  3d). The t-SNE visualisations from other LMs and different settings can be found in Appendix D.
@@
As a further analysis, we look at all the corresponding embeddings and apply k-means clustering from scikit-learn individually for backchannels/fillers in the English and Japanese data. The motivation for this analysis is based on the fact that backchannels/fillers, as functional words, often have different pragmatic functions (e.g., indicating agreement, hesitation, etc.; Petukhova and Bunt 2009). A similar idea has been discussed in the annotation work of Figueroa et al. (2022, fig. 4), who reveals that the same backchannel/filler can have several different communicative functions. In terms of our study targets, for example, the Japanese backchannel/filler ‘うん’(un, ‘yeah’) can indicate both ‘confirmation’ and ‘hesitation’. When LMs have limited knowledge of backchannels/fillers and are asked to generate the representation, models will usually give random vector values to them, which will lead to two possible clustering results: either a large $k$ value or small $k$ value. Therefore, if fine-tuning can improve the representation learning of backchannels/fillers, we should expect that the clustering effect of their embeddings will be more salient after fine-tuning. That is to say, in general we should see an increase in $k$ value, which can indicate that fine-tuned LM’s meaning representation can reflect different pragmatic functions of backchannels/fillers if $k$ was small; In contrast, if the $k$ value is initially large, it should be smaller after fine-tuning.
In order to quantitatively analyse the quality of clustering of backchannels/fillers’ embeddings before and after fine-tuning, we introduce silhouette scores Rousseeuw (1987), which are used to measure the quality of clustering. Given a range of $k$ values, which are used to perform $k$-means clustering, we calculate the corresponding silhouette coefficient s(i) for a data point $i$, which is a $n$-dim vector:
$\textbf{s(i)}=\frac{b(i)-a(i)}{\max(a(i),b(i))},$
where $a(i)$ is the average euclidean distance between the embedding of the backchannel/filler $i$ and all other embeddings in the same cluster as $i$. $b(i)$ is the minimum average euclidean distance from embedding $i$ to all embeddings in any other cluster, $1\leq s(i)\leq 1$. The silhouette score (SC) for a clustering is then measured as the average silhouette coefficient over all $n$ embeddings:
$\textbf{SC}=\frac{1}{n}\sum_{i=1}^{n}s(i)$
@@
The final result is summarised in Figure 4. In all of the three selected LMs, a general tendency we can observe is that with different fine-tuning strategies, the average silhouette score increases (with the exception of Japanese GPT-2, where the effect of fine-tuning seems not be to obvious). Detailed statistics can be found in in Tables 6 to 12 in Appendix D.
Figure 5: Evaluation metrics of generated backchannels/fillers in the English and Japanese NLG tasks. The models are evaluated before and after fine-tuning with respect to frequency, diversity, frequency-weighted perplexity, BERTScore (F1), and BLEUScore. See Tables 4 and 5 in Appendix for more details.
@@ 5 Discussion
In Section 4, the t-SNE visualization (Figure 3 reveals that fine-tuning LMs will change the representation of backchannels/fillers in a positive way. The analysis using $k$-means clustering and silhouette scores further shows that fine-tuning has beneficial effects on the representation of backchannels/fillers. Here, we further discuss the results shown in Section 4 in order to address our research questions.
RQ1: Can modern language models learn representations for backchannels/fillers through fine-tuning?
From Figure 4, as well as Tables 6 to 12, we can see that fine-tuning the pre-trained LMs with dialogue data that has contextual information on backchannels/fillers, leads to a general increase of the silhouette scores, which indicates a more salient clustering effect after fine-tuning and serves as a crucial evidence that LMs can learn the representation of backchannels/fillers with the selected fine-tuning strategies, while supplementary Figure 25 in the Appendix shows that the cluster numbers $k$ become more salient: slight increasing when $k$ was small; decreasing when $k$ was large.
@@
For both fine-tuned and non-fine-tuned LMs, a generally decreasing trend in the average silhouette score is observed with increasing context size for obtaining embeddings (exceptions are LLaMA-3 for Japanese under fine-tuning setting and Qwen-3 without fine-tuning). We interpret this result as follows: When context information is added, LMs tend to treat backchannels/fillers as functional words, whose representations would be compressed into more limited regions of the semantic space. The increased context size encodes more information from the surrounding content words to the backchannels/fillers, which “dilutes” their representations to the average level.
@@
In this paper, we targeted four types of language models: BERT and GPT-2 are smaller in size (parameters) and monolingual; LLaMA-3 8B and Qwen-3 8B are much larger and multilingual. It turns, see Figure 4 out that the average silhouette scores of clustering from LLaMA-3 embeddings are best among the four selected LM. Model size, however, does not necessarily correlate with better representation in terms of silhouette, e.g., fine-tuned BERT (which is smallest in size) has comparable results to Qwen-3.
@@
For the GPT-2-based models, we employed NTP and TTP fine-tuning strategies. The results show that both strategies help the LM learn backchannels/fillers, with no significant differences between the strategies. However, the embeddings learnt by NTP produce slightly larger silhouette scores (see Figure 4 than those learnt by TTP. Although turn-taking is considered highly relevant to the use of fillers and backchannels, our experiment shows that TTP does not obviously improve the learning of backchannel and filler representations compared to NTP. Another surprising point is that Masking for Japanese BERT model makes a significant improvement on the representation of Japanese backchannels/fillers, which is comparable to the equivalences of LLaMA-3 8B and Qwen-3-8B.
@@ 6 Further Analysis via NLG Evaluation
Besides using the silhouette score to evaluate the representation learning of backchannels/fillers in LMs, we also examine generation results to provide evidence of improved representation after fine-tuning. For evaluation, we randomly selected 20% of utterances from both the English and Japanese corpora. The generation task is defined as follows: given two turns of a dialogue in English or Japanese, both pre-trained and fine-tuned LMs are guided to continue writing the dialogue. Based on the generated content, we evaluate backchannels/fillers along five dimensions: (i) frequency of backchannels/fillers, (ii) diversity of backchannel/filler types, (iii) frequency-weighted perplexity of generated backchannels/fillers, and (iv) BERTScore (F1) and (v) BLEUScore against the ground-truth responses. The results are summarized in Figure 5, with detailed results provided in Tables 4 and 5 in Appendix.
After fine-tuning, we observe a general increase in both the frequency and diversity of generated backchannels/fillers, as well as improvements in BERTScore (except for Qwen-3 Japanese) and BLEUScore. At the same time, the frequency-weighted perplexity of backchannels/fillers decreases, offering further evidence that LMs achieve better representations of backchannels/fillers after fine-tuning. We include four illustrative examples in Figure 6 to provide additional qualitative analysis on the generated dialogue compared to ground truth. As we were concerned that fine-tuning LMs in this study might undermine the LM’s general capability of language understanding, we investigate this issue with an additional task (reported in Table 3 in the Appendix), which indicates almost no side effect.
@@ 7 Conclusion
Within the scope of this study, we consider the following limitations, which we believe can be further addressed in future work. First of all, at the beginning of the data selection for our experiment, we did consider including language resources such as the German corpora Verbmobil (VM2) (Kay, 1992) and MUNDEX (Türk et al., 2023) in order to give our results a broader linguistic basis. However, in the end we exclude the German corpora due to their comparatively small size and less formatted annotation, which, for now, leaves us with only English and Japanese data.
Secondly, the study would be more thorough with additional tests of large language models, such as Gemini and GPT-4. This was difficult given limited computing resources, where fine-tuning language models like LLaMA-3-8B was our limit. Moreover, a further step will be analysing the representation of backchannels/fillers in different hidden layers, instead of focusing on the last hidden layer of the models. There are a number of papers that examine different layers of models to answer the interpretability question (e.g., Jawahar et al., 2019; Zhao et al., 2024). We leave this as future work.
Thirdly, in this paper, we only consider how fine-tuning tasks affect representation learning of backchannels/fillers. We did not study what kind of effects different fine-tuning techniques can bring to the representation learning of backchannels/fillers. We notice that some studies propose different fine-tuning techniques. For example, surgical fine-tuning (Lee et al., 2023) selects subsets of layers to perform fine-tuning while preserving weights in other layers, which can match or outperform commonly used fine-tuning approaches.
@@
Finally, although we have performed qualitative analysis on our generated dialogues from linguistic perspective to check the use of backchannels and fillers, we currently do not have a large user study, involving human evaluation to further support our claim. We plan recruit native speakers of English and Japanese to evaluate the use of backchannels/fillers in the generated dialogues.
@@
Given the scope of this study, there do not appear to be any ethical issues. We checked the content of the selected dialogue data and made sure that there is no leakage of participants’ personal information such as name and id.

### v2

@@ 3.3 Experimental Set-Up
To illustrate the workflow, we use the backchannel uh-huh as an example. In the initial stage, we pass through all input samples $|\chi|$ and filter out all utterances that contain uh-huh. We add special token <ds> to mark and represent the backchannel (the backchannel uh-huh is transcribed as <ds> uh-huh <ds>). We then encode the corresponding utterances through the different hidden layers of a fine-tuned LM, obtaining the representation vector of uh-huh from the final hidden layer. Next, we check the dimensions of the representation matrix. Given that some of the backchannels/fillers can consist of more than one token (as uh-huh in our example), we then apply dimension reduction, simply taking the weighted average value of one dimension of the matrix so that the matrix can be levelled down to a vector representation (embedding). With the embeddings of the backchannels/fillers ‘uh-huh’, we examine its representation via clustering. For BERT and GPT-2 models used in the fine-tuning task, the dimension of the hidden layer is 768 for English GPT-2 model and 1024 for the Japanese GPT-2 model. TurnGPT is based on the GPT-2, thus shares the same dimension. A special case is LLaMA-3 8B and Qwen-3 8B as their hidden layer dimension is 4096. For computational reasons, we reduce the dimensionality of the obtained embeddings from all of the models to 100 using Principal Component Analysis (PCA; Abdi and Williams, 2010). See Appendix C for technical details of the fine-tuning (e.g., GPU run time, use of LoRA for parameter optimisation when fine-tuning the LLaMA model Hu et al., 2022).
@@
Moreover, for extracting the embeddings of backchannels/fillers after fine-tuning, we have three different context settings: (i) no context information; (ii) one context information; (iii) full context information. In the first setting, when we encounter a backchannel/filler in an utterance, only the utterance containing that backchannel/filler is fed to obtain its embedding. In the second setting, we use the previous and subsequent utterances of a backchannel/filler to build the context before obtaining its embedding. In the third setting, we combine all previous utterances when we encounter an utterance containing a backchannel/filler, using this combination as the input to obtain the embedding of that backchannel/filler. This setting is only applicable to the LLaMA-3 8B and Qwen-3 8B models since both have no input length limitation.
@@ 4 Analysis and Results
We chose the top 15 most frequent backchannels/fillers in our Japanese and English data to check how the representation of backchannels/fillers changes after fine-tuning. Among the 15 selected backchannels/fillers in each language, typical examples are: the ones indicating positive or negative feedback as the reference objects, and signals for turn-holding (in English, yes, uh, yeah, etc.; in Japanese, examples are ‘はい’ (hai, ‘yes’), ‘うん’ (un, ‘yeah’), ‘ああ’ (aa, ‘ah’), etc.).
We report the preliminary results using t-SNE visualizations (van der Maaten and Hinton, 2008). The embeddings of the 15 selected backchannels/fillers in each language are the input data. These embeddings are extracted from the last hidden layer of the LMs, before and after fine-tuning, as shown in Figure 2. We have selected t-SNE visualisations from the LLaMA-3 model with a single context setting to demonstrate how fine-tuning (in this instance, NTP) modifies the representation of backchannels/fillers within the language models (LMs). Figure 3a shows that for English data, when we obtain the embeddings of a selected backchannel/filler from the pre-trained LLaMA-3 model, the distinction between different backchannels/fillers are not clear enough. We believe that the large overlap of different data points is due to the fact that the pre-trained LMs have limited knowledge of the backchannel/filler and thus will assign random values to the encountered backchannel/filler. After fine-tuning, a clearer distinction among the embeddings of different backchannels/fillers starts to appear in the English data (see Figure 3b). The effect of fine-tuning for Japanese data is similar. After fine-tuning distinctions between different backchannels/fillers emerge as clearer boundaries among different colours appear in the t-SNE visualization (Figure 3c and  3d). The t-SNE visualisations from other LMs and different settings can be found in Appendix D.5 (Figures 10 to 19).
(a) no_ft, one_context
(b) ft, one_context (NTP)
Figure 5: Distance matrices for the top 15 Japanese backchannels/fillers in the Qwen-3 model (a) before and (b) after fine-tuning.
@@
As a further analysis, we look at all the corresponding embeddings and apply k-means clustering (from scikit-learn) individually for backchannels/fillers in the English and Japanese data. This analysis is motivated by the fact that backchannels/fillers, as functional words, often have different pragmatic functions (e.g., indicating agreement, hesitation, etc.; Petukhova and Bunt 2009). A similar idea was discussed in annotation work, which revealed that the same backchannel/filler can serve several different communicative functions (Figueroa et al., 2022, fig. 4). In terms of our study targets, for example, the Japanese backchannel/filler ‘うん’ (un, ‘yeah’) can indicate both ‘confirmation’ and ‘hesitation’. When LMs have limited knowledge of backchannels/fillers and are asked to generate their representation, models will usually give random vector values to them, which will lead to two possible clustering results: either a large or a small $k$ value. Therefore, if fine-tuning can improve the representation learning of backchannels/fillers, we should expect that the clustering effect of their embeddings will be more salient after fine-tuning. That is to say, in general we should see an increase in $k$ value, which can indicate that fine-tuned LM’s meaning representation can reflect different pragmatic functions of backchannels/fillers if $k$ was small; in contrast, if the $k$ value is initially large, it should be smaller after fine-tuning.
We first look at how our proposed approaches can distinguish the meaning representations of different backchannels/fillers. We perform $k$-means clustering and select the associated centroids (embedding values) of the backchannels/fillers before and after fine-tuning (NTP and MASK). The centroids are later averaged to produce a single representative embedding for each backchannel/filler. We select the top-15 most frequent backchannels/fillers and use confusion matrices (based on Euclidean distance) to capture the difference before and after fine-tuning.
Figure 5 shows a representative confusion matrix, where Qwen-8B is fine-tuned on Japanese dialogue data under the one-context setting (see Appendix D.7 for the other confusion matrices). In the confusion matrix, darker colours indicate greater distances and lighter colours indicate greater similarity. As can be seen, the inter-difference among the different backchannels/filler is more salient after fine-tuning. For example, the embeddings of the backchannel ‘ね’ (ne, ‘right?’, ‘hmm’, etc.) are separated from the embeddings of ‘はい’ (hai, ‘yes’, ‘mm-hmm’, etc.)
To quantitatively analyse the quality of the clustering of the embeddings of backchannels/fillers before and after fine-tuning, we use silhouette scores as a measure Rousseeuw (1987). Given a range of $k$-values, used for $k$-means clustering, we calculate the corresponding silhouette coefficient $\textbf{s}(i)$ for a data point $i$ (a backchannel/filler):
$\mathbf{s}(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))},$
where $a(i)$ is the average Euclidean distance between the embedding of the backchannel/filler $i$ and all other embeddings in the same cluster as $i$. $b(i)$ is the minimum average Euclidean distance from embedding $i$ to all embeddings in any other cluster, $1\leq\mathbf{s}(i)\leq 1$. The silhouette score (SC) for a clustering is then measured as the average silhouette coefficient over all $n$ embeddings:
$\textbf{SC}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(i),$
@@
The final result is summarised in Figure 4. In all of the four selected LMs, a general tendency we can observe is that with different fine-tuning strategies, the average silhouette score increases. Detailed statistics can be found in in Tables 7 to 14 in Appendix D.6.
Figure 6: Evaluation metrics of generated backchannels/fillers in the English and Japanese NLG tasks. The models are evaluated before and after fine-tuning with respect to frequency, diversity, frequency-weighted perplexity, BERTScore (F1), and BLEUScore (details such as metric definition and performance results are in Section D.4, Tables 4 and 5 in the Appendix.).
@@ 5 Discussion
In Section 4, the t-SNE visualization (Figure 3) reveals that fine-tuning LMs will change the representation of backchannels/fillers in a positive way. The analysis using $k$-means clustering and silhouette scores further shows that fine-tuning has beneficial effects on the representation of backchannels/fillers. Here, we further discuss these results in light of our research questions.
RQ1: Can the representations for backchannels/fillers in modern language models be learned/improved through fine-tuning?
From Figure 4 (as well as Tables 7 to 14 in the Appendix), we can see that fine-tuning the pre-trained LMs with dialogue data that contains contextual information on backchannels/fillers leads to a general increase of silhouette scores, which indicates a more salient clustering effect after fine-tuning and serves as crucial evidence that LMs can learn the representation of backchannels/fillers with the selected fine-tuning strategies. At the same time Figure 20 (Appendix) shows that the numbers of clusters $k$ become more salient: slightly increasing when $k$ was small and decreasing when $k$ was large.
@@
For both fine-tuned and not fine-tuned LMs, The average silhouette score generally exhibits an inverse relationship with the context size used for embeddings (as shown in Figure 4, exceptions are LLaMA-3 and BERT for Japanese under the fine-tuning setting, and Qwen-3 without fine-tuning). We interpret this result as follows: When context information is added, LMs tend to treat backchannels/fillers as functional words, whose representations would be compressed into more limited regions of the semantic space. The increased context size encodes more information from the surrounding content words to the backchannels/fillers, which “dilutes” their representations to the average level.
@@
In this paper, we targeted four types of language models: BERT and GPT-2 are smaller in size (parameters) and monolingual; LLaMA-3 8B and Qwen-3 8B are much larger and multilingual. It turns out (Figure 4) that the average silhouette scores of clustering from the LLaMA-3 embeddings are best among the four selected LMs. Model size, however, does not necessarily correlate with better representation in terms of silhouette scores, e.g., the results from the fine-tuned BERT model (smallest in size) are comparable to those of the Qwen-3 model.
@@
For the GPT-2-based models, we employed NTP and TTP fine-tuning strategies. The results show that both strategies help the LM learn backchannels/fillers, with no significant differences between the strategies. However, the embeddings learnt by NTP produce slightly larger silhouette scores (Figure 4 than those learnt by TTP. Although turn-taking is widely regarded as being highly relevant to the use of fillers and backchannels, our experiment shows that TTP does not significantly improve the learning of backchannel and filler representations compared to NTP. Another surprising result is that the MASK fine-tuning task for the Japanese BERT model makes significantly improves the representation of Japanese backchannels/fillers, comparable to the the results obtained with LLaMA-3 8B and Qwen-3-8B.
@@ 6 Further Analysis via NLG Evaluation
Besides using silhouette score to evaluate the representation learning of backchannels/fillers in LMs, we also examine generation results to provide evidence of improved representation after fine-tuning. For this evaluation, we randomly selected 20% of the utterances from both the English and Japanese corpora. The generation task is defined as follows: given two turns of a dialogue in English or Japanese, both pre-trained and fine-tuned (NTP) LMs are guided to continue writing the dialogue. Based on the generated content, we evaluate backchannels/fillers along five dimensions against the ground-truth responses: (i) frequency of backchannels/fillers, (ii) diversity of backchannel/filler types, (iii) frequency-weighted perplexity of generated backchannels/fillers, (iv) BERTScore (F1), and (v) BLEUScore. The results are summarized in Figure 6 (additional details are in Appendix D.4).
After fine-tuning, we observe a general increase in both the frequency and diversity of generated backchannels/fillers, as well as improvements in BERTScore (except for Qwen-3 Japanese) and BLEUScore. At the same time, the frequency-weighted perplexity of backchannels/fillers decreases, offering further evidence that LMs achieve better representations of backchannels/fillers after fine-tuning. We use six illustrative examples in Figure 9 (Appendix D.3) to provide qualitative human evaluation on the generated dialogues compared to the ground truth.
As we were concerned that fine-tuning LMs in this study might undermine a LM’s general capability of language understanding, we investigate this issue with an additional task (reported in Table 3, Appendix D.2), which indicates minor side effects.
@@ 7 Conclusion
Within the scope of this study, we consider the following limitations, which we believe can be further addressed in future work. First of all, at the beginning of the data selection for our experiment, we did consider including language resources such as the German corpora Verbmobil (VM2) (Kay et al., 1992) and MUNDEX (Türk et al., 2023) in order to give our results a broader linguistic basis. However, in the end we exclude the German corpora due to their comparatively small size and less formatted annotation, which, for now, leaves us with only English and Japanese data.
Secondly, the study would be more thorough with additional tests of large language models, such as Gemini and GPT-4. This was difficult given limited computing resources, where fine-tuning language models like LLaMA-3-8B was our limit. Moreover, a further step will be analysing the representation of backchannels/fillers in different hidden layers, instead of focusing on the last hidden layer of the models. There are a number of papers that examine different layers of models to answer the interpretability question (e.g., Jawahar et al., 2019; Zhao et al., 2024). Although we have examined the representation of backchannels/fillers in different hidden layers (results in Appendix D.1), a more systematic analysis is considered as our future work.
Thirdly, in this paper, we only consider how fine-tuning tasks affect representation learning of backchannels/fillers. We did not study what kind of effects different fine-tuning techniques can bring to the representation learning of backchannels/fillers. We notice that some studies propose different fine-tuning techniques, e.g., surgical fine-tuning (Lee et al., 2023), which selects subsets of layers to perform fine-tuning while preserving weights in other layers.
@@
Finally, we chose to conduct qualitative analysis on our generated dialogues from linguistic perspective to check the use of backchannels and fillers (details reported in Appendix D.3), which we believe adheres to the generally suggested practice when it comes to NLG evaluation: qualitative text analysis is recommended when the goal is to improve the system van der Lee et al. (2021). However, we currently do not have a large-scale user study, involving comprehensive human evaluation to further support our claim. We plan recruit native speakers of English and Japanese to evaluate the use of backchannels/fillers in the generated dialogues.
Supplementary Material
Code and data are available at https://github.com/colalao/discourse_markers (for your convenience), and as a future-proof data publication at https://doi.org/10.5281/zenodo.19473821.
@@
Given the scope of this study, there do not appear to be any ethical issues. All of the data and models used in this study are openly available or open weights. We checked the content of the selected dialogue data and made sure that there is no leakage of participants’ personal information such as name and ID. Refinement of text and experiment code were supported by ChaptGPT and GitHub Copilot.